# Volatility Spillovers in High-Dimensional Financial Systems: A Machine Learning Approach

### v1

@@ 1 Introduction
The volatility spillover literature has been dominated by GARCH-based methods since engle1982autoregressive and bollerslev1986generalized established the foundational framework. Extensions to capture asymmetric responses to positive and negative shocks include the GJR-GARCH model of glosten1993relation, which remains widely used for modeling the leverage effect in equity volatility. For multivariate settings, engle2002dynamic introduced DCC-GARCH, which models time-varying correlations parsimoniously, while engle1995multivariate developed the BEKK parameterization that allows explicit modeling of volatility transmission across assets. These multivariate GARCH specifications remain the dominant way to study volatility spillovers—appearing in hundreds of empirical applications across equity, commodity, and currency markets.
@@
. A parallel literature has developed around the spillover index of diebold2009measuring and diebold2012better, which uses variance decompositions from vector autoregressions to measure connectedness across markets. This approach has an advantage. It provides intuitive measures of directional spillovers (how much of asset A’s forecast error variance is attributable to shocks from asset B) without imposing the parametric structure of GARCH. barunik2018measuring extended this framework to the frequency domain, decomposing spillovers into short-run and long-run components. However, many Diebold-Yilmaz applications use squared returns rather than range-based or high-frequency realized volatility estimators. And the underlying VARs? Usually unrestricted, which becomes problematic as the number of assets grows.
@@
For realized volatility forecasting specifically, corsi2009simple developed the Heterogeneous Autoregressive (HAR) model, capturing the long-memory properties of volatility using a simple additive structure of daily, weekly, and monthly volatility components. The HAR model has become a workhorse for univariate volatility forecasting. But it doesn’t address cross-market spillovers. Extensions incorporating exogenous variables, like the HAR-X framework, have been developed by andersen2007roughing for jump components and applied to volatility transmission by bubak2011volatility and corsi2012discrete. audrino2016lassoing demonstrate that LASSO regularization effectively identifies sparse coefficient structures in HAR models when many potential predictors are available. However, these extensions typically focus on forecasting rather than impulse response analysis.
@@
A growing literature applies regularization methods to HAR and VAR models for spillover analysis. demirer2018estimating demonstrated the use of LASSO-VAR within the diebold2012better spillover framework. With this approach they were able to estimate the high-dimensional network of 150 global banks. For commodity markets, yang2021volatility apply LASSO-VAR to 25 commodity futures, finding that energy commodities serve as primary spillover transmitters. In the HAR literature, ma2019forecasting show that Elastic Net and LASSO outperform individual HAR models for oil price volatility—with Elastic Net performing better due to its ability to handle correlated predictors. audrino2016lassoing provide theoretical and empirical support for applying LASSO to HAR models, showing that regularization effectively recovers sparse structures without sacrificing forecast accuracy. ding2021forecasting find that multivariate LASSO models outperform univariate HAR at longer forecast horizons and introduce ordered LASSO to respect time series structure. An important finding from this literature? Regularization enables network identification even when it doesn’t substantially improve forecasting. buccheri2021multivariate document that cross-market spillovers in multivariate HAR models have long-term forecasting power, but LASSO lag selection doesn’t outperform parsimonious HAR specifications. That result foreshadows our own finding: HAR slightly outperforms our multivariate models while the multivariate structure remains essential for understanding spillover dynamics.
@@
Our approach bridges these literatures by developing a hybrid HAR-ElasticNet framework that addresses two distinct econometric challenges in high-dimensional volatility modeling: capturing realistic persistence dynamics and identifying sparse cross-market spillover networks. Standard unrestricted VAR estimation becomes infeasible when the number of variables approaches the sample size. With six markets and three HAR horizons per market, our model includes 108 potential parameters. Our two-step procedure first estimates univariate Heterogeneous Autoregressive (HAR) models via OLS for each market’s own-volatility dynamics, preserving the characteristic high persistence ($\approx 0.99$ at daily frequency) documented in the realized volatility literature (corsi2009simple). We then apply ElasticNet regularization to the cross-market terms, setting approximately 70–75 percent of spillover parameters to zero. This sparsity is economically sensible—while a handful of central markets exhibit significant cross-linkages, most market pairs have negligible direct volatility transmission. By separating own-dynamics from cross-market selection, our approach combines the strengths of HAR models (realistic persistence) with regularization methods (sparse network identification), yielding impulse responses that reveal volatility spillover structure.
@@
Second, consistent with buccheri2021multivariate, the parsimonious HAR model and our hybrid HAR-ElasticNet achieve identical out-of-sample RMSE (0.0044). This equivalence reflects the hybrid model’s sparsity. For forecasting purposes, the hybrid model performs similarly to the univariate HAR. This demonstrates that cross-market information contributes negligibly to predictive accuracy while revealing economically meaningful network structure.
@@
Our contribution is both methodological and empirical. We extend the regularized HAR literature (buccheri2021multivariate; ma2019forecasting) to a multi-asset-class setting spanning commodities, equities, and treasuries, and combine spillover network identification with Joint IRF analysis for policy-relevant insights into cross-market volatility transmission. To the best of our knowledge, this is the first application of the JIRF concept to a multivariate volatility modeling exercise.
@@ 2.2 Realized Volatility Estimation
We compute realized volatility using the Yang-Zhang estimator (yang2000drift), which is considered the most efficient range-based volatility estimator. Using a 30-day rolling window:
@@
where $\sigma_{o}^{2}$ is the overnight variance (variance of $\log(O_{t}/C_{t-1})$), $\sigma_{oc}^{2}$ is the open-to-close variance (variance of $\log(C_{t}/O_{t})$), and $\sigma_{RS}^{2}$ is the rolling mean of the Rogers-Satchell (rogers1991estimating) daily variance:
@@
The Yang-Zhang estimator is more efficient than Garman-Klass (garman1980estimation) or Parkinson (parkinson1980extreme) estimators because it’s robust to opening jumps (overnight gaps) while still using all available daily price information. yang2000drift show it has the minimum variance among unbiased range-based estimators for assets that exhibit opening jumps—which is common in futures markets.
@@ 3.1 The HAR Model for Own-Volatility Dynamics
Realized volatility exhibits two stylized facts that inform our modeling approach: high persistence (autocorrelations that decay slowly over many lags) and apparent long memory (corsi2009simple). The Heterogeneous Autoregressive (HAR) model captures these properties parsimoniously by aggregating volatility at daily, weekly, and monthly horizons:
@@ 3.4 HAR Benchmark
Our primary benchmark is the univariate HAR model of corsi2009simple. For each asset $i$:
@@ 3.6 Joint Impulse Response Functions
Traditional IRFs measure the response to a one-unit shock in a single variable. Joint Impulse Response Functions (JIRF), developed by wiesen2023joint, generalize this framework to simultaneous shocks across multiple variables. This extension is particularly valuable for studying financial market stress, where events such as Federal Reserve announcements or geopolitical crises typically affect multiple markets at once rather than one in isolation. The JIRF methodology allows us to trace out the dynamic effects of such multi-market shocks while properly accounting for the joint structure of the initial impulse.
@@
Joint Shock Construction. When shocking multiple variables simultaneously (e.g., both ES and NQ in an equity shock), we must specify how the shocks are correlated. Following wiesen2023joint, we use generalized impulse response functions (GIRFs ) in the spirit of pesaran1998generalized, which condition on the estimated residual covariance structure rather than imposing an arbitrary ordering as in Cholesky decomposition. Specifically, when we shock variable $i$ by one standard deviation, we simultaneously adjust other shocked variables $j\in\mathcal{S}$ according to their conditional expectation given the covariance matrix $\hat{\Sigma}_{u}$ of the hybrid HAR-ElasticNet residuals. For the equity shock, this means that the ES and NQ shocks are correlated according to their estimated contemporaneous correlation (approximately 0.85 in our sample), reflecting the empirical reality that equity volatility shocks do not occur independently across indices. This approach avoids the ordering sensitivity of Cholesky-based identification while preserving the joint structure of multi-market volatility events. This approach loses causal interpretation, but implements shocks that are empirically consistent with observed correlations.
@@ 4.1 Forecast Performance
Consistent with buccheri2021multivariate, parsimonious specifications without cross-market spillovers achieve equivalent out-of-sample forecast performance to multivariate models. Table 3 reports out-of-sample performance using an 80/20 train-test split: models are estimated on the first 4,559 observations (May 2002 to July 2020) and evaluated on the remaining 1,140 observations (July 2020 to January 2025) via rolling one-step-ahead forecasts. [^2] See the online supplement for MAE and MAPE estimates of forecast performance comparisons. They demonstrate identical results for both models. Both the univariate HAR benchmark and the hybrid HAR-ElasticNet achieve identical average RMSE of 0.0044. The hybrid model’s extreme sparsity—only 8% of cross-market coefficients are nonzero—effectively reduces it to a univariate HAR specification for forecasting purposes, confirming that cross-market spillovers contribute negligibly to predictive accuracy, despite their economic significance for network structure (Sections 4.2–4.5).
This equivalence between forecast accuracy and model complexity is central to the spillover literature. The HAR model exploits the well-documented long-memory properties of realized volatility through its three-component autoregressive structure (corsi2009simple), efficiently capturing own-market persistence that dominates any predictive signal from cross-market dynamics. The hybrid HAR-ElasticNet model’s aggressive regularization on cross-market terms yields identical forecasts to univariate HAR while preserving the ability to identify economically meaningful spillover channels. This demonstrates that volatility transmission networks, though critical for understanding systemic risk propagation, provide minimal incremental forecasting power beyond own-market persistence.
@@
Notes: Out-of-sample RMSE for one-step-ahead forecasts. Test period: July 2020 – January 2025 (1,140 observations). Training period: May 2002 – July 2020 (4,559 observations). Bold indicates best performance per row. HAR is the Heterogeneous Autoregressive model of corsi2009simple. The hybrid HAR-ElasticNet model achieves identical forecast performance to univariate HAR, reflecting its extreme sparsity (only 8% nonzero cross-market coefficients). This demonstrates that cross-market spillovers, while economically significant for network structure (Sections 4.2–4.5), contribute negligibly to predictive accuracy.
@@ 5 Conclusion
Consistent with buccheri2021multivariate, the univariate HAR model and our hybrid specification achieve identical out-of-sample RMSE (0.0044), reflecting the hybrid model’s extreme sparsity: with only 8% nonzero cross-market coefficients, it effectively reduces to univariate HAR for forecasting purposes. This demonstrates that cross-market information adds negligible signal for point prediction.

### v2

# Two-Step Regularized HARX to Measure Volatility Spillovers in Multi-Dimensional Systems

@@ 1 Introduction
The volatility spillover literature has been dominated by GARCH-based methods since Engle (1982) and Bollerslev (1986) established the foundational framework. Extensions to capture asymmetric responses to positive and negative shocks include the GJR-GARCH model of Glosten et al. (1993), which remains widely used for modeling the leverage effect in equity volatility. For multivariate settings, Engle (2002) introduced DCC-GARCH, which models time-varying correlations parsimoniously, while Engle and Kroner (1995) developed the BEKK parameterization that allows explicit modeling of volatility transmission across assets. These multivariate GARCH specifications remain the dominant way to study volatility spillovers—appearing in hundreds of empirical applications across equity, commodity, and currency markets.
@@
. A parallel literature has developed around the spillover index of Diebold and Yilmaz (2009) and Diebold and Yilmaz (2012), which uses variance decompositions from vector autoregressions to measure connectedness across markets. This approach has an advantage. It provides intuitive measures of directional spillovers (how much of asset A’s forecast error variance is attributable to shocks from asset B) without imposing the parametric structure of GARCH. Baruňík and Křehlík (2018) extended this framework to the frequency domain, decomposing spillovers into short-run and long-run components. However, many Diebold-Yilmaz applications use squared returns rather than range-based or high-frequency realized volatility estimators. And the underlying VARs? Usually unrestricted, which becomes problematic as the number of assets grows.
@@
For realized volatility forecasting specifically, Corsi (2009) developed the Heterogeneous Autoregressive (HAR) model, capturing the long-memory properties of volatility using a simple additive structure of daily, weekly, and monthly volatility components. The HAR model has become a workhorse for univariate volatility forecasting. But it doesn’t address cross-market spillovers. Extensions incorporating exogenous variables, like the HAR-X framework, have been developed by Andersen et al. (2007) for jump components and applied to volatility transmission by Bubák et al. (2011) and Corsi and Renò (2012). Audrino and Knaus (2016) demonstrate that LASSO regularization effectively identifies sparse coefficient structures in HAR models when many potential predictors are available. However, these extensions typically focus on forecasting rather than impulse response analysis.
@@
A growing literature applies regularization methods to HAR and VAR models for spillover analysis. Demirer et al. (2018) demonstrated the use of LASSO-VAR within the Diebold and Yilmaz (2012) spillover framework. With this approach they were able to estimate the high-dimensional network of 150 global banks. For commodity markets, Yang and Zhou (2021) apply LASSO-VAR to 25 commodity futures, finding that energy commodities serve as primary spillover transmitters. In the HAR literature, Ma et al. (2018) show that Elastic Net and LASSO outperform individual HAR models for oil price volatility—with Elastic Net performing better due to its ability to handle correlated predictors. Audrino and Knaus (2016) provide theoretical and empirical support for applying LASSO to HAR models, showing that regularization effectively recovers sparse structures without sacrificing forecast accuracy. Ding et al. (2021) find that multivariate LASSO models outperform univariate HAR at longer forecast horizons and introduce ordered LASSO to respect time series structure. An important finding from this literature? Regularization enables network identification even when it doesn’t substantially improve forecasting. Buccheri et al. (2021) document that cross-market spillovers in multivariate HAR models have long-term forecasting power, but LASSO lag selection doesn’t outperform parsimonious HAR specifications. That result foreshadows our own finding: HAR slightly outperforms our multivariate models while the multivariate structure remains essential for understanding spillover dynamics.
@@
Our approach bridges these literatures by developing a hybrid HAR-ElasticNet framework that addresses two distinct econometric challenges in high-dimensional volatility modeling: capturing realistic persistence dynamics and identifying sparse cross-market spillover networks. Standard unrestricted VAR estimation becomes infeasible when the number of variables approaches the sample size. With six markets and three HAR horizons per market, our model includes 108 potential parameters. Our two-step procedure first estimates univariate Heterogeneous Autoregressive (HAR) models via OLS for each market’s own-volatility dynamics, preserving the characteristic high persistence ($\approx 0.99$ at daily frequency) documented in the realized volatility literature (Corsi, 2009). We then apply ElasticNet regularization to the cross-market terms, setting approximately 70–75 percent of spillover parameters to zero. This sparsity is economically sensible—while a handful of central markets exhibit significant cross-linkages, most market pairs have negligible direct volatility transmission. By separating own-dynamics from cross-market selection, our approach combines the strengths of HAR models (realistic persistence) with regularization methods (sparse network identification), yielding impulse responses that reveal volatility spillover structure.
@@
Second, consistent with Buccheri et al. (2021), the parsimonious HAR model and our hybrid HAR-ElasticNet achieve identical out-of-sample RMSE (0.0044). This equivalence reflects the hybrid model’s sparsity. For forecasting purposes, the hybrid model performs similarly to the univariate HAR. This demonstrates that cross-market information contributes negligibly to predictive accuracy while revealing economically meaningful network structure.
@@
Our contribution is both methodological and empirical. We extend the regularized HAR literature (Buccheri et al., 2021; Ma et al., 2018) to a multi-asset-class setting spanning commodities, equities, and treasuries, and combine spillover network identification with Joint IRF analysis for policy-relevant insights into cross-market volatility transmission. To the best of our knowledge, this is the first application of the JIRF concept to a multivariate volatility modeling exercise.
@@ 2.2 Realized Volatility Estimation
We compute realized volatility using the Yang-Zhang estimator (Yang and Zhang, 2000), which is considered the most efficient range-based volatility estimator. Using a 30-day rolling window:
@@
where $\sigma_{o}^{2}$ is the overnight variance (variance of $\log(O_{t}/C_{t-1})$), $\sigma_{oc}^{2}$ is the open-to-close variance (variance of $\log(C_{t}/O_{t})$), and $\sigma_{RS}^{2}$ is the rolling mean of the Rogers-Satchell (Rogers and Satchell, 1991) daily variance:
@@
The Yang-Zhang estimator is more efficient than Garman-Klass (Garman and Klass, 1980) or Parkinson (Parkinson, 1980) estimators because it’s robust to opening jumps (overnight gaps) while still using all available daily price information. Yang and Zhang (2000) show it has the minimum variance among unbiased range-based estimators for assets that exhibit opening jumps—which is common in futures markets.
@@ 3.1 The HAR Model for Own-Volatility Dynamics
Realized volatility exhibits two stylized facts that inform our modeling approach: high persistence (autocorrelations that decay slowly over many lags) and apparent long memory (Corsi, 2009). The Heterogeneous Autoregressive (HAR) model captures these properties parsimoniously by aggregating volatility at daily, weekly, and monthly horizons:
@@ 3.4 HAR Benchmark
Our primary benchmark is the univariate HAR model of Corsi (2009). For each asset $i$:
@@ 3.6 Joint Impulse Response Functions
Traditional IRFs measure the response to a one-unit shock in a single variable. Joint Impulse Response Functions (JIRF), developed by Wiesen and Beaumont (2024), generalize this framework to simultaneous shocks across multiple variables. This extension is particularly valuable for studying financial market stress, where events such as Federal Reserve announcements or geopolitical crises typically affect multiple markets at once rather than one in isolation. The JIRF methodology allows us to trace out the dynamic effects of such multi-market shocks while properly accounting for the joint structure of the initial impulse.
@@
Joint Shock Construction. When shocking multiple variables simultaneously (e.g., both ES and NQ in an equity shock), we must specify how the shocks are correlated. Following Wiesen and Beaumont (2024), we use generalized impulse response functions (GIRFs ) in the spirit of Pesaran and Shin (1998), which condition on the estimated residual covariance structure rather than imposing an arbitrary ordering as in Cholesky decomposition. Specifically, when we shock variable $i$ by one standard deviation, we simultaneously adjust other shocked variables $j\in\mathcal{S}$ according to their conditional expectation given the covariance matrix $\hat{\Sigma}_{u}$ of the hybrid HAR-ElasticNet residuals. For the equity shock, this means that the ES and NQ shocks are correlated according to their estimated contemporaneous correlation (approximately 0.85 in our sample), reflecting the empirical reality that equity volatility shocks do not occur independently across indices. This approach avoids the ordering sensitivity of Cholesky-based identification while preserving the joint structure of multi-market volatility events. This approach loses causal interpretation, but implements shocks that are empirically consistent with observed correlations.
@@ 4.1 Forecast Performance
Consistent with Buccheri et al. (2021), parsimonious specifications without cross-market spillovers achieve equivalent out-of-sample forecast performance to multivariate models. Table 3 reports out-of-sample performance using an 80/20 train-test split: models are estimated on the first 4,559 observations (May 2002 to July 2020) and evaluated on the remaining 1,140 observations (July 2020 to January 2025) via rolling one-step-ahead forecasts. [^2] See the online supplement for MAE and MAPE estimates of forecast performance comparisons. They demonstrate identical results for both models. Both the univariate HAR benchmark and the hybrid HAR-ElasticNet achieve identical average RMSE of 0.0044. The hybrid model’s extreme sparsity—only 8% of cross-market coefficients are nonzero—effectively reduces it to a univariate HAR specification for forecasting purposes, confirming that cross-market spillovers contribute negligibly to predictive accuracy, despite their economic significance for network structure (Sections 4.2–4.5).
This equivalence between forecast accuracy and model complexity is central to the spillover literature. The HAR model exploits the well-documented long-memory properties of realized volatility through its three-component autoregressive structure (Corsi, 2009), efficiently capturing own-market persistence that dominates any predictive signal from cross-market dynamics. The hybrid HAR-ElasticNet model’s aggressive regularization on cross-market terms yields identical forecasts to univariate HAR while preserving the ability to identify economically meaningful spillover channels. This demonstrates that volatility transmission networks, though critical for understanding systemic risk propagation, provide minimal incremental forecasting power beyond own-market persistence.
@@
Notes: Out-of-sample RMSE for one-step-ahead forecasts. Test period: July 2020 – January 2025 (1,140 observations). Training period: May 2002 – July 2020 (4,559 observations). Bold indicates best performance per row. HAR is the Heterogeneous Autoregressive model of Corsi (2009). The hybrid HAR-ElasticNet model achieves identical forecast performance to univariate HAR, reflecting its extreme sparsity (only 8% nonzero cross-market coefficients). This demonstrates that cross-market spillovers, while economically significant for network structure (Sections 4.2–4.5), contribute negligibly to predictive accuracy.
@@ 5 Conclusion
Consistent with Buccheri et al. (2021), the univariate HAR model and our hybrid specification achieve identical out-of-sample RMSE (0.0044), reflecting the hybrid model’s extreme sparsity: with only 8% nonzero cross-market coefficients, it effectively reduces to univariate HAR for forecasting purposes. This demonstrates that cross-market information adds negligible signal for point prediction.